OpenStudy (anonymous):

The frequency of vibrations of a piano string varies directly as the square root of the tension on the string and inversely as the length of the string. The middle A string has a frequency of 440 vibrations per second. Find the frequency of a string that has 1.25 times as much tension and is 1.2 times as long.

OpenStudy (anonymous):

were you able to derive an equation based on the information supplied ?

OpenStudy (anonymous):

No, I am terrible at word problems

OpenStudy (anonymous):

let f = frequency we know it is directly proportional (varies directly) as the square root of tension. we can call that \[\sqrt{T}\] Frequency also varies inversely (varies indirectly) as the length of the string. we can call that \[L\] \[f = \frac{ k \sqrt{T} }{ L }\] where k is the constant of proportionality. will this information help ?

OpenStudy (anonymous):

Could you work it with me

OpenStudy (anonymous):

try deriving an equation for k using the fact that the frequency is 440 vibrations per second.

OpenStudy (anonymous):

So would it be 440=k(square root 1.25) divided by 1.2

OpenStudy (anonymous):

\[f_{A} = \frac{ k \sqrt{T_{A}} }{ L_{A} } = 440 \rightarrow k = \frac{ 440 L_{A} }{ \sqrt{T_{A}} }\] The new string has a tension 1.25 as much as Ta and length 1.2 x as much, so our new string as a frequency . . . \[f_{S} = \frac{ k \sqrt{T _{S}}}{ L _{S} }\] We know Ts = 1.25 Ta, and Ls = 1.2 La, and k is given above. so finally: \[f_{S} = \frac{ 440 L_{A} }{ \sqrt{T_{A}} } * \frac{ \sqrt{1.25 T_{A}} }{ 1.2 L_{A} }\] cancel the units and obtain the desired result

OpenStudy (anonymous):

so would my answer be fs =410

OpenStudy (anonymous):

that sounds reasonable

OpenStudy (anonymous):

and the calculator agrees

OpenStudy (anonymous):

is it 410 vibrations per second

OpenStudy (anonymous):

yep. i get 409.946... so 410

OpenStudy (anonymous):

Thank you